ammonia gas is passed into water yielding a solution of density 0.93 g/cm3 and containing 18.6% NH3 by weight the mass of NH3 per cc of the solution is

Grade:11

1 Answers

Arun
4 years ago
Density = Mass/Volume 
So, Mass = Density * Volume
 Density = 0.93g/cc 
Mass of 1 cc of solution ={0.93 * 1 }= 0.93g 
 1 cc of solution =0.93g
 Solution contains 18.6% ammonia by weight.
 So,1 cc of solution =18.6% of 0.93g ammonia= (18.6/100) * 0.93 = 0.173g 
So, this comes the mass of ammonia per cc of solution which is 0.173g.